OpenStudy (aaronandyson):

A point P on x-axis is equidistant from the point (7,6) and (-4,3).Find the co-ordinates of the point P.

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (aaronandyson):

@amistre64

OpenStudy (michele_laino):

If I call with \((x,y)\) the coordinate of point \(P\), then I can write this: \[\sqrt {{{\left( {x - 7} \right)}^2} + {{\left( {y - 6} \right)}^2}} = \sqrt {{{\left( {x - \left( { - 4} \right)} \right)}^2} + {{\left( {y - 3} \right)}^2}} \]

OpenStudy (amistre64):

a line between 2 points, is full of points that are equidistant

OpenStudy (aaronandyson):

P is on x-axis so y should be 0.

OpenStudy (michele_laino):

correct! if we take the square of both sides, after a simplification I get: \[{\left( {x - 7} \right)^2} + {\left( {0 - 6} \right)^2} = {\left( {x + 4} \right)^2} + {\left( {0 - 3} \right)^2}\]

OpenStudy (amistre64):

pfft, thought it said P is a point ... well i spose 'on the x axis' narrows it down lol

OpenStudy (aaronandyson):

x = 30/11?

OpenStudy (michele_laino):

correct!

OpenStudy (aaronandyson):

Thanks.